OpenStudy (buttermequeasy):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (radar):

slope is the coefficient of x and this case it is 3. Parallel lines will have the same slope (3) 3 = (y-(-4))/(x-2) cross multiplying 3x-6 = y+4 y=3x-10 looks like the first one